OpenStudy (ineedhelpfast12):

write an equation in standard form of the line with the given characteristics. Slope: 10; passes through (6,2)

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

(x1,y1)=(6,2) & m=10 substitute the values in the equation and you're done :)

OpenStudy (ineedhelpfast12):

wait so 6-6*2?

OpenStudy (anonymous):

y-2=10*(x-6)

OpenStudy (ineedhelpfast12):

oh, so m would be the slope, x1 the x-intercept, and y1 the y-intercept

OpenStudy (anonymous):

yup

OpenStudy (ineedhelpfast12):

danke